Solve

x-1/3-4 = 3x-8/12

Thank You!!!

Answer by stanbon(75887)

[(x-1)/3]-4 = (3x-8)/12
Notice how you should have posted your problem.
Multiply through by 12 to get:
4(x-1)-48 = 3x-8
4x-4-48 = 3x-8
x-52=-8
x=44
Cheers,
Stan H.